# Modeling Developer Burnout with GenAI Adoption

### v1

@@ 5. Impact of AI Adoption on Burnout (RQ1)
To address RQ1, we followed a concurrent embedded mixed-methods approach (Greene, 2007; Morse, 2016) (as shown in Figure 1). We first leveraged Partial Least Squares–Structural Equation Modeling (PLS–SEM) (Ringle et al., 2015) to analyze the survey data to evaluate hypotheses H1 to H3 through steps: - . We then conducted qualitative analysis on open-ended question responses to capture developers’ perspectives on how AI adoption reshapes demands and resources ().
@@ 5.1. Method
PLS-SEM analysis, a multivariate analysis technique, is our analysis of choice as it allows simultaneous testing of multiple relationships among constructs, while accounting for measurement error in modeling latent constructs (e.g., Burnout) (Trinkenreich et al., 2024b). It is also increasingly being used in empirical SE to investigate complex phenomena (Trinkenreich et al., 2024a; Choudhuri et al., 2024). The JD-R theoretical foundation, as explained in Section 3, serves as the guiding lens for our analysis and includes the steps outlined in Figure 1.
@@
Measurement Model. We modeled Burnout as a construct measured by four reflective indicators from our survey items. These indicators capture different yet complementary facets of burnout: workload manageability (Burn1), emotional and physical exhaustion (Burn2), cynicism/disengagement (Burn3), and reduced efficacy/job insecurity (Burn4). In a reflective measurement model, the indicators covary—e.g., when developers are emotionally or physically exhausted, they are less effective—and the construct (Burnout) is inferred from their shared variance.
@@
Job Demands was modeled as a formative higher-order construct (HOC) comprising two lower-order constructs (LOCs)– Organizational Pressure and Workload. We modeled these LOCs as formative indicators of Job Demands (HOC) for both theoretical and methodological reasons. First, consistent with the JD–R Literature, job demands represent a set of heterogeneous stressors that collectively increase strain, but need not covary, that is, they are not interchangeable reflections of a single latent trait. For example, a high workload does not necessarily imply high organizational pressure; yet, both independently contribute to burnout. Second, formative modeling allows us to account for the fact that removing one LOC (e.g., workload) would fundamentally alter the Job Demands construct, rather than simply reducing measurement reliability. This treatment aligns with prior applications of the JD–R model, which have modeled different types of demands as complementary contributors to stress and burnout (Schaufeli and Bakker, 2004).
@@
The LOC Organizational Pressure modeled as a latent construct measured by four reflective indicators (Org1–Org4), which captured structural and managerial expectations tied to AI adoption (e.g., “Management expects me to use AI”). Similarly, Workload was measured by four reflective indicators (Work1-Work4) that capture the perception of intensified task requirements (e.g., “I often have to do more work than I can handle well”).
@@
Job Resources was modeled as a formative HOC composed of two reflectively measured LOCs: Autonomy and Learning Resources. These two constructs are modeled as formative indicators of Job Resources (HOC), as they neither covary nor are they interchangeable (e.g., having autonomy does not imply getting AI training resources, and vice versa). This is in line with prior research, which has modeled job resources formatively when combining distinct categories such as social support, feedback, and autonomy (Schaufeli and Bakker, 2004). Autonomy, a latent construct, was measured using three reflective indicators from our survey items assessing developers’ influence over how they carry out their tasks (Auto1), their authority to handle responsibilities (Auto2), and their ability to make important decisions about their work (Auto3). Learning Resources was a single-item measure capturing whether developers had access to organizational support for acquiring AI-related knowledge and skills.
@@
AI-perception was modeled as a control variable and measured through a single item capturing developers’ overall attitudes toward AI adoption, which may influence developer strain (Davis, 1989; Venkatesh et al., 2012). For the complete mapping of survey items to indicators of constructs, see Section 2 in the supplementary document (Anonym, 2025).
@@
Sample Size and Power Analysis. To ensure the adequacy of our sample size, we conducted a power analysis analysis using the G*Power tool (Faul et al., 2009). We have three predictors of burnout in our model (Job Demands, Job Resources, and AI Perceptions as a control). Assuming a medium effect size ($f^{2}=.15$), significance level ($\alpha=.05$), and desired statistical power ($1-\beta=.95$), the analysis indicated a minimum required sample size of 119. Our sample size of 442 far exceeds this requirement.
@@
Exploratory Factor Analysis (EFA) was needed because, although the JD–R constructs have been validated in prior literature (Scholze and Hecker, 2024; Roșca et al., 2021; Dorta-Afonso et al., 2023), we had to adapt the survey items to our context of developer workflows and AI adoption. EFA allowed us to verify that the items clustered in ways that matched our intended constructs and that the underlying factor structure was consistent with our JD–R model (Fabrigar et al., 1999; Worthington and Whittaker, 2006). All EFAs were performed using JASP (JASP Team, 2023).
@@
Following established guidance (Costello and Osborne, 2005; Worthington and Whittaker, 2006), we first confirmed that the data met the necessary assumptions for factor analysis (1) sampling adequacy (Kaiser–Meyer–Olkin $\geq.60$) (Hair, 2009), and (2) sufficient intercorrelations among items (a significant Bartlett’s test of sphericity) (Bartlett, 1950). We then evaluated items against multiple criteria: (3) factor loadings ($\geq.50$ preferred, .30.30.50.50 considered marginal) (Hair, 2009; Hair et al., 2019), (4) cross-loadings (<.30<.30) (Hair et al., 2019), (5) uniqueness (<.60<.60) (Hair, 2009), and (6) parallel analysis (keeping only the factors with eigenvalues larger than the corresponding average random eigenvalues) and inspection of the scree plot (keeping factors before the curve levels off) for final determination of factor retention (Howard, 2016). See Tables 2-4 in the supplementary document for the test results (Anonym, 2025)
@@
The results supported a two-factor structure for Job Demands (Organizational Pressure and Workload), a one-factor structure for Job Resources (Autonomy), and a one-factor structure for Burnout. Because Learning Resources was measured with a single item, it was not included in the EFA (Hair, 2009; Worthington and Whittaker, 2006). Work1, Work2, Burn2, and Burn4 items had elevated uniqueness (>.60>.60), but were retained because their factor loadings exceeded the recommended threshold of .50, which is considered a strong indicator of construct relevance in social science research (Costello and Osborne, 2005; Fabrigar et al., 1999).
@@
PLS-SEM analysis. We then conducted PLS-SEM analysis using SmartPLS 4 (Ringle et al., 2015). PLS-SEM involves two steps: () evaluating the measurement model to empirically assess the relationships between the construct and its constituent indicators, and () evaluating the structural model to test the hypothesized relationships among constructs. These results are reported in Section 5.2.
@@
⑥Composite Assessment. To better understand developers’ experiences with AI adoption, we qualitatively analyzed participants’ responses ($N=221$) to the open-ended survey question: “If you have any final thoughts or experiences on how adopting AI tools has impacted your work, please share them below.” We used LOCs (e.g., Organizational Pressure, Workload, Autonomy, Learning Resources) and the control variable (AI-Perception) from our model as our codeset. Two authors independently coded 20% of the responses, and compared their outputs using the Jaccard index, achieving 90% agreement on inter-rater reliability (IRR) (Landis and Koch, 1977). The remaining responses were then divided between the two authors and coded individually using the stable codebook. See codebook in supplementary document Table 15 (Anonym, 2025).
5.2. Results
@@
Measurement Model Assessment. We used a two-stage evaluation to assess the (1) reflective measurement model and (2) formative measurement model as per (Becker et al., 2012; Sarstedt et al., 2019).
@@
Stage 1: Reflective Measurement Model Evaluation. We evaluated the reflective constructs (LOCs: Organizational Pressure, Workload, Autonomy, Learning Resources) and the endogenous construct (Burnout) by using tests for: (i) convergent validity, (ii) internal consistency reliability, (iii) discriminant validity, and (iv) collinearity among reflective indicators (Hair et al., 2019; Hair Jr et al., 2020).
@@
(i) Convergent validity examines whether indicators intended to measure the same construct share sufficient variance (Kock, 2014). We assessed this using (a) outer loadings and (b) Average Variance Extracted (AVE) (Hair et al., 2019). (a) Outer loadings represent the bivariate relation between each indicator and its latent construct; values $\geq.70$ shows that the indicator shares at least 50% of its variance with the construct. Indicators with outer loadings between 0.400.400.700.70 can be removed when doing so improves construct-level reliability, or AVE, or if they are theoretically essential (Hair Jr et al., 2020). In our model, as shown in Figure 2, Work2 (0.610), Org4 (0.574), and Burn4 (0.66) fell within this range, but were retained because removing them did not improve construct-level reliability. All other indicators exceed the recommended threshold of 0.70. (b) AVE summarizes the proportion of common variance across indicators and should exceed $0.50$, indicating that the latent construct explains at least half of the variance in its indicators (Hair et al., 2019). Table 2 shows that the AVEs meet the criteria.
@@
(ii) Internal consistency reliability determines whether indicators consistently reflect their underlying construct and is assessed using Cronbach’s $\alpha$ and Composite Reliability (CR: $\rho_{a}$, $\rho_{c}$) tests (Russo and Stol, 2021). The acceptable range for both these values is between $0.70$ and $0.9$ (Hair et al., 2019). Table 2 shows all LOCs and Burnout is in the acceptable range.
@@
$\mathrm{Cronbach’s}$ $\alpha$ $\mathrm{CR}(\rho_{a})$ $\mathrm{CR}(\rho_{c})$ $\mathrm{AVE}$
@@
Cronbach’s $\alpha$ tends to underestimate reliability, whereas composite reliability ($\mathrm{CR}:\rho_{c}$) tends to overestimate it. The true reliability typically lies between these two estimates and is effectively captured by $\mathrm{CR}(\rho_{a})$ (Russo and Stol, 2021). Note: Learning Resources was measured with a single item and is therefore not included in this table, as reliability and AVE are not applicable for single-item construct.
(iii) Discriminant validity determines whether each construct is empirically distinct from the others and includes three tests: (a) the Heterotrait–Monotrait (HTMT) ratio of correlations, where values below .90 indicate adequate discriminant validity; (b) indicator cross-loadings, where each item should load higher on its intended construct than on others; and (c) the Fornell–Larcker criterion, where the square root of each construct’s AVE should be greater than its correlations with other constructs (Henseler et al., 2015; Hair et al., 2019). All constructs meet these criteria and are reported in Tables 7-9 in the supplementary document (Anonym, 2025).
(iv) Collinearity check of reflective indicators ensures that indicators did not exhibit multicollinearity among themselves. We used the Variance Inflation Factor (VIF), where values below 3 to 5 are considered acceptable (Hair et al., 2019), and all indicators in our model satisfy this criterion (see Table 10 in the supplementary doc (Anonym, 2025)).
Stage 2: Formative Measurement Model Evaluation. We then evaluated the formative constructs (HOCs) following the assessment steps (Becker et al., 2012; Sarstedt et al., 2019): (i) the significance of outer weights, (ii) outer loadings, and (iii) collinearity among the LOCs for each HOC.
@@
(ii) Outer loadings assess whether each LOC contributes meaningfully to its HOC by reflecting their bivariate correlation. Values above $0.50$ are generally considered acceptable evidence of relevance, even when outer weights are marginal (Sarstedt et al., 2019). In our case, the outer loading of Organizational Pressure is >0.5>0.5 and therefore is retained, although it has marginally significant outer weights.
@@
(iii) Test for collinearity among LOCs ensures that each HOC’s constituent LOCs do not exhibit multicollinearity. We used the VIF, and all values are well below the recommended cutoffs (below 3). See Tables 11-13 for Stage 2 analysis in the supplementary (Anonym, 2025).
@@
Structural Model Evaluation. We evaluated the structural model by (i) assessing collinearity among predictor constructs, (ii) testing the hypotheses, and (iii) examining the explanatory power of the model. Figure 2 presents the model.
@@
(i) Collinearity among predictor constructs was assessed, and it was found that predictor constructs in the structural model (Job Demands, Job Resources, and AI Perceptions) do not exhibit multicollinearity by calculating VIFs. All VIF values are well below the threshold of 3 (see Table 14 in supplementary (Anonym, 2025)).
@@
Figure 2. Outer loadings and path coefficients (p ¡ 0.05 indicated by a full line). Higher order constructs (e.g. Job-Demands) are represented by “double circles” and have paths to their corresponding lower order constructs (e.g. Workload); a dashed rectangle represents “Control Variable” (e.g. AI-Perception).
@@
(ii) Path coefficients were estimated to identify the direct effects of one construct on another in the structural model. Path coefficients were assessed using nonparametric bootstrapping with 5,000 subsamples, which provides empirical standard errors, $p$-values, and confidence intervals for inference (Hair et al., 2019). Table 3 reports the path coefficients ($\beta$), standard deviations (SD), 95% confidence intervals, and $p$-values.
@@
We found empirical support for hypotheses H1 and H2. As shown in Figure 2, Job Demands (i.e., organizational pressures and workload intensification) are positively associated with burnout ($\beta=0.398$, $p<.001$). In contrast, Job Resources related to GenAI adoption (i.e., autonomy and learning resources) are negatively associated with burnout ($\beta=-0.360$, $p<.001$). For the HOC constructs of Job Demands, Workload has a stronger contribution, while for Job Resources, Autonomy has a stronger contribution. Hypothesis H3 also has empirical support. AI perception is negatively associated with burnout ($\beta=-0.246$, $p<.001$), indicating that more favorable AI perceptions correspond to lower burnout.
@@
(iii) Explanatory Power of Burnout was assessed using the coefficient of determination ($R^{2}$). The $R^{2}$ value for Burnout is 0.398. (Studies in organizational and social science contexts note that $R^{2}$ values around .30–.40 represent moderate explanatory power, given the complexity of human and organizational behavior (Cohen, 2013; Chin, 1998).)
@@
We assessed the model’s external validity by assessing its predictive relevance using Stone–Geisser’s $Q^{2}$ via blindfolding in SmartPLS. Blindfolding iteratively omits and predicts portions of the data (using an omission distance $D$) and computes $Q^{2}$ from prediction errors, where values above 0 indicate predictive relevance (Stone, 1974; Hair Jr et al., 2021). In our case, the $Q^{2}$ value for Burnout is 0.373, indicating a large predictive relevance.
@@
We assessed the model fit using the standardized root mean square residual (SRMR) (Russo and Stol, 2021), with a value less than 0.08 (conservative) or 0.10 (more lenient) considered a good fit. Our models’ SRMR meets the threshold.
@@
Table 3. Standardized path coefficients (B), standard devia- tions (SD), confidence intervals (95% CI), p-values ($p$).
@@
Composite Assessment. To complement the quantitative findings, our qualitative analysis captures developers’ perspectives on how AI adoption reshapes job demands and resources.

### v2

@@ 5. Impact of AI Adoption on Burnout (RQ1)
Overview of the analysis process for addressing RQ1: 1. Measurement model definition 2. Sample size and power analysis 3. Exploratory factor analysis (EFA) 4. Measurement model assessment 5. Structural model evaluation (PLS-SEM) 6. Composite assessment
To address RQ1, we followed a concurrent embedded mixed-methods approach (greene2007mixed; morse2016mixed) (as shown in Figure 1). We first leveraged Partial Least Squares–Structural Equation Modeling (PLS–SEM) (ringle2015structural) to analyze the survey data to evaluate hypotheses H1 to H3 through steps: - . We then conducted qualitative analysis on open-ended question responses to capture developers’ perspectives on how AI adoption reshapes demands and resources ().
@@ 5.1. Method
PLS-SEM analysis, a multivariate analysis technique, is our analysis of choice as it allows simultaneous testing of multiple relationships among constructs, while accounting for measurement error in modeling latent constructs (e.g., Burnout) (trinkenreich2024predicting). It is also increasingly being used in empirical SE to investigate complex phenomena (trinkenreich2025investigating; choudhuri2024far). The JD–R theoretical foundation, as explained in Section 3, serves as the guiding lens for our analysis and includes the steps outlined in Figure 1.
@@
①Measurement Model. We modeled Burnout as a construct measured by four reflective indicators from our survey items. These indicators capture different yet complementary facets of burnout: workload manageability (Burn1), emotional and physical exhaustion (Burn2), cynicism/disengagement (Burn3), and reduced efficacy/job insecurity (Burn4). In a reflective measurement model, the indicators covary—e.g., when developers are emotionally or physically exhausted, they are less effective—and the construct (Burnout) is inferred from their shared variance.
@@
Job Demands was modeled as a formative higher-order construct (HOC) comprising two lower-order constructs (LOCs)– Organizational Pressure and Workload. We modeled these LOCs as formative indicators of Job Demands (HOC) for both theoretical and methodological reasons. First, consistent with the JD–R Literature, job demands represent a set of heterogeneous stressors that collectively increase strain, but need not covary, that is, they are not interchangeable reflections of a single latent trait. For example, a high workload does not necessarily imply high organizational pressure; yet, both independently contribute to burnout. Second, formative modeling allows us to account for the fact that removing one LOC (e.g., workload) would fundamentally alter the Job Demands construct, rather than simply reducing measurement reliability. This treatment aligns with prior applications of the JD–R model, which have modeled different types of demands as complementary contributors to stress and burnout (schaufeli2004job).
@@
The LOC Organizational Pressure modeled as a latent construct measured by four reflective indicators (Org1-Org4), which captured structural and managerial expectations tied to AI adoption (e.g., “Management expects me to use AI”). Similarly, Workload was measured by four reflective indicators (Work1-Work4) that capture the perception of intensified task requirements (e.g., “I often have to do more work than I can handle well”).
@@
Job Resources was modeled as a formative HOC composed of two reflectively measured LOCs– Autonomy and Learning Resources. These two constructs are modeled as formative indicators of Job Resources (HOC), as they neither covary nor are they interchangeable (e.g., having autonomy does not imply getting AI training resources, and vice versa). This is in line with prior research, which has modeled job resources formatively when combining distinct categories such as social support, feedback, and autonomy (schaufeli2004job). Autonomy, a latent construct, was measured using three reflective indicators from our survey items assessing developers’ influence over how they carry out their tasks (Auto1), their authority to handle responsibilities (Auto2), and their ability to make important decisions about their work (Auto3). Learning Resources was a single-item (Res) measure capturing whether developers had access to organizational support for acquiring AI-related knowledge and skills.
@@
AI-perception was modeled as a control variable and measured through a single item capturing developers’ overall attitudes toward AI adoption, which may influence developer strain (davis1989perceived; venkatesh2012consumer). For the complete mapping of survey items to indicators of constructs, see Section 2 in the supplementary document (supply).
@@
②Sample Size and Power Analysis. To ensure the adequacy of our sample size, we conducted a power analysis using the G*Power tool (faul2009statistical). We have three predictors of burnout in our model (Job Demands, Job Resources, and AI Perceptions as a control). Assuming a medium effect size ($f^{2}=.15$), significance level ($\alpha=.05$), and desired statistical power ($1-\beta=.95$), the analysis indicated a minimum required sample size of 119. Our sample size of 442 far exceeds this requirement.
@@
③Exploratory Factor Analysis (EFA) was needed because, although the JD–R constructs have been validated in prior literature (scholze2024job; rosca2021job; dorta2023s), we had to adapt the survey items to our context of developer workflows and AI adoption. EFA allowed us to verify that the items clustered in ways that matched our intended constructs and that the underlying factor structure was consistent with our JD–R model (fabrigar1999evaluating; worthington2006scale). All EFAs were performed using JASP (jasp2023).
@@
Following established guidance (costello2005best; worthington2006scale), we first confirmed that the data met the necessary assumptions for factor analysis (1) sampling adequacy (Kaiser–Meyer–Olkin $\geq.60$) (hair2009multivariate), and (2) sufficient intercorrelations among items (a significant Bartlett’s test of sphericity) (bartlett1950tests). We then evaluated items against multiple criteria: (3) factor loadings ($\geq.50$ preferred, .30.30.50.50 considered marginal) (hair2009multivariate; hair2019use), (4) cross-loadings (<.30<.30) (hair2019use), (5) uniqueness (<.60<.60) (hair2009multivariate), and (6) parallel analysis (keeping only the factors with eigenvalues larger than the corresponding average random eigenvalues) and inspection of the scree plot (keeping factors before the curve levels off) for final determination of factor retention (howard2016review). See Tables 2-4 in the supplementary document for the test results (supply).
@@
The results supported a two-factor structure for Job Demands (Organizational Pressure and Workload), a one-factor structure for Job Resources (Autonomy), and a one-factor structure for Burnout. Because Learning Resources was measured with a single item, it was not included in the EFA (hair2009multivariate; worthington2006scale). Work1, Work2, Burn2, and Burn4 items had elevated uniqueness (>.60>.60), but were retained because their factor loadings exceeded the recommended threshold of .50, which is considered a strong indicator of construct relevance in social science research (costello2005best; fabrigar1999evaluating).
@@
PLS-SEM analysis. We then conducted PLS-SEM analysis using SmartPLS 4 (ringle2015structural). PLS-SEM involves two steps: () evaluating the measurement model to empirically assess the relationships between the construct and its constituent indicators, and () evaluating the structural model to test the hypothesized relationships among constructs. These results are reported in Section 5.2.
@@
⑥Composite Assessment. To better understand developers’ experiences with AI adoption, we qualitatively analyzed participants’ responses ($N=221$) to the open-ended survey question: “If you have any final thoughts or experiences on how adopting AI tools has impacted your work, please share them below.” We used LOCs (e.g., Organizational Pressure, Workload, Autonomy, Learning Resources) and the control variable (AI-Perception) from our model as our codeset. Two authors independently coded 20% of the responses, and compared their outputs using the Jaccard index, achieving 90% agreement on inter-rater reliability (IRR) (landis1977measurement). The remaining responses were then divided between the two authors and coded individually using the stable codebook. See codebook in the supplementary document Table 15 (supply).
5.2. Results for RQ1
@@
④Measurement Model Assessment. We used a two-stage evaluation to assess the (1) reflective measurement model and (2) formative measurement model as per becker2012hierarchical; sarstedt2019specify.
@@
Stage 1: Reflective Measurement Model Evaluation. We evaluated the reflective constructs (LOCs: Organizational Pressure, Workload, Autonomy, Learning Resources) and the endogenous construct (Burnout) by using tests for: (i) convergent validity, (ii) internal consistency reliability, (iii) discriminant validity, and (iv) collinearity among reflective indicators (hair2019use; hair2020assessing).
@@
(i) Convergent validity examines whether indicators intended to measure the same construct share sufficient variance (kock2014advanced). We assessed this using (a) outer loadings and (b) Average Variance Extracted (AVE) (hair2019use). (a) Outer loadings represent the bivariate relation between each indicator and its latent construct; values $\geq.70$ shows that the indicator shares at least 50% of its variance with the construct. Indicators with outer loadings between 0.400.400.700.70 can be removed when doing so improves construct-level reliability, or AVE, or if they are theoretically essential (hair2020assessing). In our model, as shown in Figure 2, Work2 (0.61), Org4 (0.574), and Burn4 (0.66) fell within this range, but were retained because removing them did not improve construct-level reliability. All other indicators exceed the recommended threshold of 0.70. (b) AVE summarizes the proportion of common variance across indicators and should exceed $0.50$, indicating that the latent construct explains at least half of the variance in its indicators (hair2019use). Table 2 shows that the AVEs meet the criteria.
@@
(ii) Internal consistency reliability determines whether indicators consistently reflect their underlying construct and is assessed using Cronbach’s $\alpha$ and Composite Reliability (CR: $\rho_{a}$, $\rho_{c}$) tests (russo2021pls). The acceptable range for both these values is between $0.70$ and $0.9$ (hair2019use). Table 2 shows all LOCs and Burnout is in the acceptable range.
@@
$\mathrm{Cronbach^{\prime}s}$ $\alpha$ $\mathrm{CR}(\rho_{a})$ $\mathrm{CR}(\rho_{c})$ $\mathrm{AVE}$
@@
Cronbach’s $\alpha$ tends to underestimate reliability, whereas composite reliability ($\mathrm{CR}:\rho_{c}$) tends to overestimate it. The true reliability typically lies between these two estimates and is effectively captured by $\mathrm{CR}(\rho_{a})$ (russo2021pls). Note: Learning Resources was measured with a single item and is therefore not included in this table, as reliability and AVE are not applicable for single-item construct.
Internal Consistency Reliability and Convergent Validity
(iii) Discriminant validity determines whether each construct is empirically distinct from the others and includes three tests: (a) the Heterotrait-Monotrait (HTMT) ratio of correlations, where values below .90 indicate adequate discriminant validity; (b) indicator cross-loadings, where each item should load higher on its intended construct than on others; and (c) the Fornell-Larcker criterion, where the square root of each construct’s AVE should be greater than its correlations with other constructs (henseler2015new; hair2019use). All constructs meet these criteria and are reported in Tables 7-9 in the supplementary document (supply).
(iv) Collinearity check of reflective indicators ensures that indicators did not exhibit multicollinearity among themselves. We used the Variance Inflation Factor (VIF), where values below 3 to 5 are considered acceptable (hair2019use), and all indicators in our model satisfy this criterion (see Table 10 in the supplementary document (supply)).
Stage 2: Formative Measurement Model Evaluation. We then evaluated the formative constructs (HOCs) following the assessment steps (becker2012hierarchical; sarstedt2019specify): (i) the significance of outer weights, (ii) outer loadings, and (iii) collinearity among the LOCs for each HOC.
@@
(ii) Outer loadings assess whether each LOC contributes meaningfully to its HOC by reflecting their bivariate correlation. Values above $0.50$ are generally considered acceptable evidence of relevance, even when outer weights are marginal (sarstedt2019specify). In our case, the outer loading of Organizational Pressure is >0.5>0.5 and therefore is retained, although it has marginally significant outer weights.
@@
(iii) Test for collinearity among LOCs ensures that each HOC’s constituent LOCs do not exhibit multicollinearity. We used the VIF, and all values are well below the recommended cutoffs (below 3). See Tables 11-13 for Stage 2 analysis in the supplementary (supply).
@@
⑤Structural Model Evaluation. We evaluated the structural model by (i) assessing collinearity among predictor constructs, (ii) testing the hypotheses, and (iii) examining the explanatory power of the model. Figure 2 presents the model.
@@
(i) Collinearity among predictor constructs was assessed, and it was found that predictor constructs in the structural model (Job Demands, Job Resources, and AI Perceptions) do not exhibit multicollinearity by calculating VIFs. All VIF values are well below the threshold of 3 (see Table 14 in the supplementary document(supply)).
@@
Figure 2. Outer loadings and path coefficients (p ¡ 0.05 indicated by a full line; $N=442$). Higher order constructs (e.g. Job-Demands) are represented by “double circles” and have paths to their corresponding lower order constructs (e.g. Workload); a dashed rectangle represents “Control Variable” (e.g. AI-Perception).
Outer loadings and path coefficients (p ¡ 0.05 indicated by a full line). Higher order constructs (e.g. Job-Demands) are represented by “double circles” and have paths to their corresponding lower order constructs (e.g. Workload); a dashed rectangle represents “Control Variable” (e.g. AI-Perception).
@@
(ii) Path coefficients were estimated to identify the direct effects of one construct on another in the structural model. Path coefficients were assessed using nonparametric bootstrapping with 5,000 subsamples, which provides empirical standard errors, $p$-values, and confidence intervals for inference (hair2019use). Table 3 reports the path coefficients ($\beta$), standard deviations (SD), 95% confidence intervals, and $p$-values.
@@
We found empirical support for hypotheses H1 and H2. As shown in Figure 2, Job Demands (i.e., organizational pressures and workload intensification) are positively associated with burnout ($\beta=0.398$, $p<.001$). In contrast, Job Resources related to GenAI adoption (i.e., autonomy and learning resources) are negatively associated with burnout ($\beta=-0.360$, $p<.001$). For the LOCs of Job Demands, Workload has a stronger contribution, while for Job Resources, Autonomy has a stronger contribution. Hypothesis H3 also has empirical support. AI perception is negatively associated with burnout ($\beta=-0.246$, $p<.001$), indicating that more favorable AI perceptions correspond to lower burnout.
@@
(iii) Explanatory Power of Burnout was assessed using the coefficient of determination ($R^{2}$). The $R^{2}$ value for Burnout is 0.398 (Studies in organizational and social science contexts note that $R^{2}$ values around .30.30.40.40 represent moderate explanatory power, given the complexity of human and organizational behavior (cohen2013statistical; chin1998partial)).
@@
We assessed the model’s external validity by assessing its predictive relevance using Stone-Geisser’s $Q^{2}$ via blindfolding in SmartPLS. Blindfolding iteratively omits and predicts portions of the data (using an omission distance $D$) and computes $Q^{2}$ from prediction errors, where values above 0 indicate predictive relevance (stone1974cross; hair2021partial). In our case, the $Q^{2}$ value for Burnout is 0.373, indicating a large predictive relevance.
@@
We assessed the model fit using the standardized root mean square residual (SRMR) (russo2021pls), with a value less than 0.08 (conservative) or 0.10 (more lenient) considered a good fit. Our models’ SRMR meets the threshold.
@@
Table 3. Standardized path coefficients (B), standard deviations (SD), confidence intervals (95% CI), p-values ($p$).
@@
Standardized path coefficients, standard deviations, confidence intervals (95% CI), p-values.
⑥Composite Assessment. To complement the quantitative findings, our qualitative analysis captures developers’ perspectives on how AI adoption reshapes job demands and resources.